OpenStudy (anonymous):

. Solve. P = 2l + 2w + 2k for k ?? show work!!

OpenStudy (anonymous):

answer it mathematicalpro!!

OpenStudy (anonymous):

\[2(l+w+k)=P\] \[l+w+k=P/2\] \[k=(P/2)-l-w\]

OpenStudy (anonymous):

is this right??

OpenStudy (anonymous):

or \[k=(P-2l-2w)/2\]